# Multi-Agent Retrieval-Augmented Framework for Evidence Based Counterspeech Against Health Misinformation

### v1

@@ Abstract
Large language models (LLMs) incorporated with Retrieval-Augmented Generation (RAG) have demonstrated powerful capabilities in generating counterspeech against misinformation. However, current studies rely on limited evidence and offer less control over final outputs. To address these challenges, we propose a Multi-agent Retrieval-Augmented Framework to generate counterspeech against health misinformation, incorporating multiple LLMs to optimize knowledge retrieval, evidence enhancement, and response refinement. Our approach integrates both static and dynamic evidence, ensuring that the generated counterspeech is relevant, well-grounded, and up-to-date. Our method outperforms baseline approaches in politeness, relevance, informativeness, and factual accuracy, demonstrating its effectiveness in generating high-quality counterspeech. To further validate our approach, we conduct ablation studies to verify the necessity of each component in our framework. Furthermore, human evaluations reveal that refinement significantly enhances counterspeech quality and obtains human preference.
@@ 1 Introduction
The rapid spread of misinformation on social media has become a significant concern, particularly in health-related topics. Health misinformation about disease prevention and untested treatments spreads widely, confusing the public and undermining trust in medical institutions (Chou et al., 2021; Cinelli et al., 2020). Consequently, individuals hesitate to follow public health guidelines and some even resort to harmful self-medication practices(Pennycook et al., 2020), underscoring the emergence of mitigating health misinformation. Recent studies prove that social media users can help with misinformation debunking (Chuai et al., 2024; Caulfield et al., 2020), who can rapidly respond, share factual information and correct misinformation, extending the impact of debunking efforts beyond official channels. However, they often struggle to access reliable evidence and are limited by the large volume and complexity of health misinformation (Mourali & Drake, 2022). Therefore, empowering users to effectively counter misinformation is crucial (He et al., 2023; Swire-Thompson et al., 2020).
@@
Large language models (LLMs) are powerful in response generation and are widely used in mitigating misinformation (Proma et al., 2025; Chen & Shu, 2024; Saha et al., 2024; Hong et al., 2024). However, LLMs are not always reliable and can produce hallucinations, prompting researchers to integrate external knowledge to counter misinformation(Bondarenko & Viehweger, 2024). Retrieval-augmented generation (RAG) has been employed to incorporate external knowledge, facilitating the generation of evidence-based counterspeech(Ni et al., 2025; Yue et al., 2024; Liu et al., 2024). Various studies typically rely on static knowledge (Chung et al., 2021; Guu et al., 2020), which refers to verified knowledge that is relatively stable and does not frequently change(Shrinivasan & Razniewski, 2022). For example, the U.S. Surgeon General’s Advisory on Building a Healthy Information Environment  [^1] https://www.ncbi.nlm.nih.gov presents established public health principles and guidelines for addressing health misinformation. Recently, dynamic knowledge has been increasingly incorporated into response generation(Yeginbergen et al., 2025; Komeili et al., 2021; Nakano et al., 2021). This kind of knowledge usually refers to information that is continually updated and responsive to evolving situations. For instance, the U.S. Department of Health and Human Services (HHS) has announced that it will produce 4.8 million doses of the H5N1 avian flu vaccine to prepare for a possible pandemic. This development has been reported by reliable sources such as CIDRAP [^2] https://www.cidrap.umn.edu/, CBS News [^3] https://www.cbsnews.com/, and AJMC [^4] https://www.ajmc.com/. Although recent studies have incorporated various forms of knowledge, most existing methods utilize only one type of knowledge: either the static or the dynamic, ignoring the potential benefits of interacting with both.
@@
Incorporating both static and dynamic knowledge is essential in mitigating health misinformation. Firstly, they are complementary to each other: Although the static knowledge is reliable, it has the limitation that it can not quickly adapt to emerging health misinformation or recent updates. In contrast, dynamic evidence can adapt rapidly to new contexts, but may lack nuance or rigorous validation, potentially weakening the effectiveness of counterspeech(Braun et al., 2024). Therefore, integrating both forms of knowledge allows static knowledge, such as established guidelines, to help filter out low-quality information, while dynamic knowledge ensures the latest development(Charlotte et al., 2024; Nezafat & Samet, 2024). Secondly, incorporating both types of evidence helps build public trust. Individuals are more likely to follow advice relying on longstanding, peer-reviewed research while simultaneously staying alert to contemporary changes(Yeginbergen et al., 2025; Drolsbach et al., 2024).
@@
In our study, we design a multi-source knowledge framework that encompasses both static and dynamic evidence. The static knowledge base provides reliable, focused information, built through careful manual curation. Meanwhile, the dynamic knowledge base continuously integrates real-time updates using a web search API, ensuring that the framework remains responsive to evolving health information. However, incorporating large volumes of evidence may lead to information overload, resulting in responses less accessible to users. To address this, we design a refined module to post-edit responses, ensuring they are clear, concise, and user-friendly(Madaan et al., 2023; Deng et al., 2023). In our study, we propose a Multi-Agent Retrieval-Augmented Framework that integrates various components to achieve effective counterspeech generation. Multiple agents are designed to perform various tasks, including retrieving evidence from static and dynamic knowledge bases, summarizing evidence, generating responses, and refining them for clarity and accessibility.
@@
  • Multi-Agent Retrieval-Augmented Framework
    We propose a multi-agent pipeline that integrates a designed retrieval and generation module to realize effective counterspeech generation. Our method outperforms across all evaluation dimensions (Politeness (0.88), and Informativeness (0.78) while maintaining strong Relevance (0.70) and Factual Accuracy (0.86)).
  • Evidence Enhancement and Refined Response Generation
    Our system integrates static retrieval and dynamic retrieval to gather reliable and updated evidence, ensuring robust support for countering health misinformation. Additionally, our refinement agent makes counterspeech more accessible, leading to higher human preference in our evaluation.
@@ 2.1 Evidence-based Misinformation Mitigation
Recently, several studies have leveraged LLMs to produce contextually relevant and effective responses by utilizing their own knowledge base (Mou et al., 2024; Lodovico Molina et al., 2024). However, due to the limitation of LLMs’ knowledge base, they may produce hallucinations and fabricated information (Huang et al., 2025; Rawte et al., 2023). Retrieval-augmented generation (RAG) is proposed to introduce external evidence for producing counterspeech (Ni et al., 2025; Yue et al., 2024; Leekha et al., 2024), improving the factual accuracy of the content generated. This method primarily depends on its knowledge base to provide factual information, but it does not account for health-related changes over time (Ouyang et al., 2025). Some studies combine an LLM with online web searches to identify evidence-based misinformation (Tian et al., 2024; Li et al., 2024a; Dey et al., 2024), which will mitigate the limitations of static knowledge base. These systems enhance evidence retrieval by incorporating information from diverse online sources (Shi et al., 2025). However, web searches introduce new challenges like retrieving low-quality or misleading sources. To address these issues, our study integrates both static and dynamic retrieved evidence to ensure the availability of relevant, real-time information for countering misinformation.
@@ 2.2 LLM Agents for Social Response Generation
Integrating LLMs across multiple agents has dramatically improved response generation (Zhang et al., ; Zhao et al., 2024). Multi-agent LLM systems feature several specialized agents working together to tackle complex tasks, utilizing LLMs to boost the quality and efficiency of interactions (Yang et al., 2025; Guo et al., 2024; Cruz, 2024). Many Studies utilize multiple LLM-based multi-agent architectures to improve information retrieval and response generation (Salve et al., 2024), tackling tasks that require long-context understanding (Zhang et al., ), and enhancing recommendation quality and user engagement (Fang et al., 2024). For example, Clarke et al. (2022) implemented a framework where agents specialize in different tasks but coordinate to deliver cohesive responses, utilizing a central management system. These multi-agent approaches demonstrate the effectiveness of breaking down complex response generation into modular, specialized components. Inspired by this, our Multi-agent Retrieval-Augmented Framework employs sequentially organized agents, each with distinct capabilities, to retrieve, filter, summarize, generate, and refine evidence-based counterspeech.
Figure 1: Overview of our proposed Multi-Agent Counterspeech Generation Framework for addressing health misinformation. First, the Static Retrieval Agent gathers information from a local database, while the Dynamic Retrieval Agent fetches real-time evidence online. Next, the Summarization Agent filters and condenses the information for clarity. The Counterspeech Agent then creates a response, and the Refinement Agent improves to ensure it is clear, respectful, and well-supported. This process ensures effective, evidence-based counterspeech.
@@ 3.1 Generation
The counterspeech generation process involves leveraging LLMs to create effective, informative, and polite responses that counter health misinformation. The primary objective is to generate responses that not only provide factual corrections but also address misinformation in a respectful and engaging manner. Our proposed multi-agent system uses specialized agents working in sequence to retrieve, filter, and process evidence, optimizing the RAG pipeline.
Information Retrieval To incorporate static and dynamic evidence, two separate agents are employed to retrieve evidence: one is a Static Retrieval Agent and another one is a Dynamic Retrieval Agent. These Agents are collecting information based on the misinformation posts. The Static Retrieval Agent conducts an inquiry within a structured local knowledge base by employing a dual approach that integrates vector-based semantic search with keyword-based exact matching (Yuan et al., 2024; Sawarkar et al., 2024), ensuring that only valid medical knowledge will be included in this process. In parallel, the Dynamic Retrieval Agent performs real-time web searches using the DuckDuckGo API, introducing the latest health misinformation corrections from authoritative sources. Two sources of knowledge are combined into the final evidence.
@@ 3.2 Evaluation
We employed a comprehensive evaluation framework based on four key criteria: politeness, relevance, informativeness, and factual accuracy. These criteria ensure that the responses are engaging, respectful, factually sound, and effective in addressing health misinformation.
@@
(b) Relevance: Relevant responses ensure the response directly addresses the specific misinformation claim (Yue et al., 2024; He et al., 2023). We assess relevance using the BM25 relevance scoring algorithm (Robertson et al., 2009), which measures how well the generated response aligns with the original misinformation content.
(c) Informativeness: Informativeness reflects the amount of useful, relevant, and evidence-based content shared in the response. A highly informative counterspeech response not only corrects misinformation but also provides actionable insights and explanations that improve user understanding and support informed decision-making(Chung et al., 2024). To measure informativeness, we employ a LLM as an evaluator (Li et al., 2024b) with detailed instructions, assessing if the response provides valuable information.
(d) Factual Accuracy: Factual accuracy demonstrates the reliability and trustworthiness of the generated response by ensuring that the information is correct and supported by evidence(Zhou et al., 2024). In counterspeech Generation, presenting scientifically accurate facts can effectively correct misinformation and maintaining user trust (Yue et al., 2024). We employ LLM evaluation (Li et al., 2024b) to check how accurate the generated answers are by comparing them to the gathered evidence. The model rates how accurate a response is by checking if it correctly presents scientific facts and does not include any false information.
@@ 4.1 Dataset
We utilize PRW API  [^5] https://praw.readthedocs.io/ to collect health misinformation from Reddit, focusing on topics related to the coronavirus disease (COVID-19), influenza (Flu), and human immunodeficiency virus (HIV), as these topics have been the center of conspiracy theories and misinformation in recent times. We collect Reddit posts and comments containing health-related keywords (e.g., ”vaccines,” ”COVID-19,” ”alternative medicine”, etc. See details in Appendix A.1.2)). Finally, we obtain about 4,968 posts and 17,223 comments from high-engagement subreddits (e.g., r/health or r/science, etc. See details in Appendix A.1.1).
After collecting the Reddit dataset, we employ human annotators to identify and filter health misinformation (Song et al., 2025b). Considering the annotation cost, we only extract 1000 samples for annotation. Five annotators with backgrounds in information science were recruited for annotating. To ensure consistency and reliability in annotations, all annotators were provided with comprehensive and detailed annotation guidelines Appendix A.2.1. We finally obtained 330 posts containing health misinformation.
To further enrich our dataset, we developed an automatic classifier trained on the annotated data, enabling the efficient collection of additional health misinformation posts from unannotated data. We employ the RoBERTa-large model (Liu et al., 2019), trained on our manually annotated dataset, which achieves a moderate accuracy with an F1 score of 0.76. We apply it to the remaining dataset, identifying 831 additional posts containing health misinformation. To validate the reliability of the classifier, we extract 100 samples to conduct human validation (See details in Appendix A.2.2). The final dataset contained 1,161 posts labeled as health misinformation.

### v2

# Multi-Agent Retrieval-Augmented Framework for Evidence-Based Counterspeech Against Health Misinformation

@@ Abstract
Large language models (LLMs) incorporated with Retrieval-Augmented Generation (RAG) have demonstrated powerful capabilities in generating counterspeech against misinformation. However, current studies rely on limited evidence and offer less control over final outputs. To address these challenges, we propose a Multi-agent Retrieval-Augmented Framework to generate counterspeech against health misinformation, incorporating multiple LLMs to optimize knowledge retrieval, evidence enhancement, and response refinement. Our approach integrates both static and dynamic evidence, ensuring that the generated counterspeech is relevant, well-grounded, and up-to-date. Our method outperforms baseline approaches in politeness, relevance, informativeness, and factual accuracy, demonstrating its effectiveness in generating high-quality counterspeech. To further validate our approach, we conduct ablation studies to verify the necessity of each component in our framework. Furthermore, cross evaluations show that our system generalizes well across diverse health misinformation topics and datasets. And human evaluations reveal that refinement significantly enhances counterspeech quality and obtains human preference.
@@ 1 Introduction
The rapid spread of misinformation on social media has become a significant concern, particularly in health-related topics. Health misinformation about disease prevention and untested treatments spreads widely, confusing the public and undermining trust in medical institutions (Chou et al., 2021; Cinelli et al., 2020). Consequently, individuals hesitate to follow public health guidelines, and some even resort to harmful self-medication practices (Pennycook et al., 2020), underscoring the emergence of mitigating health misinformation. Recent studies prove that social media users can help with misinformation debunking (Chuai et al., 2024; Caulfield et al., 2020), who can rapidly respond, share factual information and correct misinformation, extending the impact of debunking efforts beyond official channels. However, they often struggle to access reliable evidence and are limited by the large volume and complexity of health misinformation (Mourali & Drake, 2022). Therefore, empowering users to effectively counter misinformation is crucial (He et al., 2023; Swire-Thompson et al., 2020).
@@
Large language models (LLMs) are powerful in response generation and are widely used in mitigating misinformation (Proma et al., 2025; Chen & Shu, 2024; Saha et al., 2024; Hong et al., 2024). However, LLMs are not always reliable and can produce hallucinations, prompting researchers to integrate external knowledge to counter misinformation (Bondarenko & Viehweger, 2024). Retrieval-Augmented Generation (RAG) has been employed to incorporate external knowledge, facilitating the generation of evidence-based counterspeech (Ni et al., 2025; Yue et al., 2024; Liu et al., 2024). Various studies typically rely on static knowledge (Chung et al., 2021; Guu et al., 2020), which refers to verified knowledge that is relatively stable and does not frequently change (Shrinivasan & Razniewski, 2022). For example, the U.S. Surgeon General’s Advisory on Building a Healthy Information Environment  [^1] https://www.ncbi.nlm.nih.gov/ presents established public health principles and guidelines for addressing health misinformation. Recently, dynamic knowledge has been increasingly incorporated into response generation (Yeginbergen et al., 2025; Komeili et al., 2022; Nakano et al., 2021). This kind of knowledge usually refers to information that is continually updated and responsive to evolving situations. For instance, ”The U.S. Department of Health and Human Services (HHS) has announced that it will produce 4.8 million doses of the H5N1 avian flu vaccine to prepare for a possible pandemic.” This development has been reported by reliable sources such as CIDRAP [^2] https://www.cidrap.umn.edu/, CBS News [^3] https://www.cbsnews.com/, and AJMC [^4] https://www.ajmc.com/. Although recent studies have incorporated various forms of knowledge, most existing methods utilize only one type of knowledge: either static or dynamic, ignoring the potential benefits of interacting with both.
@@
Incorporating both static and dynamic knowledge is essential in mitigating health misinformation. Firstly, they are complementary to each other. Although the static knowledge is reliable, it has the limitation that it can not quickly adapt to emerging health misinformation or recent updates. In contrast, dynamic evidence can adapt rapidly to new contexts, but may lack nuance or rigorous validation, potentially weakening the effectiveness of counterspeech (Braun et al., 2024). Therefore, integrating both forms of knowledge allows static knowledge, such as established guidelines, to help filter out low-quality information, while dynamic knowledge ensures the latest development (Nezafat & Samet, 2024). Secondly, incorporating both types of evidence helps build public trust. Individuals are more likely to follow advice relying on longstanding, peer-reviewed research while simultaneously staying alert to contemporary changes (Yeginbergen et al., 2025; Drolsbach et al., 2024).
@@
In our study, we design a multi-source knowledge framework that encompasses both static and dynamic evidence. The static knowledge base provides reliable, focused information, built through careful manual curation. Meanwhile, the dynamic knowledge base continuously integrates real-time updates using a web search API, ensuring that the framework remains responsive to evolving health information. However, incorporating large volumes of evidence may lead to information overload, resulting in responses that are less accessible to users. To address this, we design a refined module to post-edit responses, ensuring they are clear, concise, and user-friendly (Madaan et al., 2023; Deng et al., 2023). In our study, we propose a Multi-Agent Retrieval-Augmented Framework that integrates various components to achieve effective counterspeech generation. Multiple agents are designed to perform various tasks, including retrieving evidence from static and dynamic knowledge bases, summarizing evidence, generating responses, and refining them for clarity and accessibility.
@@
  • Multi-Agent Retrieval-Augmented Framework Specialized for Counterspeech Generation Against Health Misinformation
    This pipeline presents a modular multi-agent framework tailored for generating effective counterspeech in response to health misinformation. It integrates specialized sub-agents responsible for: (1) static and dynamic retrieval, (2) evidence summarization, (3) counterspeech generation, and (4) iterative response refinement. Each agent functions autonomously yet collaboratively within a scalable architecture, enhancing the quality of the generated counterspeech.
  • Integration of Static and Dynamic Knowledge
    The framework combines both static (e.g., curated medical guidelines) and dynamic (e.g., real-time web search) evidence to generate counterspeech, enabling responses that are both authoritative and up-to-date. To our knowledge, this combination has not been explored previously, especially in the context of generating counterspeech to health misinformation. Our experiments demonstrate the importance of incorporating both static and dynamic knowledge.
  • Ablation Studies with Prompting Strategy Comparison
    Comprehensive ablation studies evaluate the contribution of each agent in the pipeline. The system is tested under different prompting strategies (e.g., Guided Prompting vs. Chain-of-Thought), demonstrating that each agent remains essential regardless of the prompting method used. Such testing has not been implemented or reported in previous research on counterspeech generation.
  • A Curated Dataset on Diverse Health Misinformation Topics
    A new dataset is introduced, featuring health misinformation posts related to COVID-19, influenza, and HIV, sourced from Reddit and annotated through a combination of classifier-assisted annotation and expert labeling. This dataset is valuable for health misinformation research and is publicly available on GitHub  [^5] https://github.com/AnirbanSahaAnik/health-misinformation-reddit-dataset/.
@@ 2.1 Evidence-based Misinformation Mitigation
Recently, several studies have leveraged LLMs to produce contextually relevant and effective responses by utilizing their own knowledge base (Mou et al., 2024; Lodovico Molina et al., 2024). However, due to the limitation of LLMs’ knowledge base, they produced hallucinations and fabricated information (Huang et al., 2025; Rawte et al., 2023). Retrieval-augmented generation (RAG) was proposed to introduce external evidence for producing counterspeech (Ni et al., 2025; Yue et al., 2024; Leekha et al., 2024), improving the factual accuracy of the content generated. This method primarily depended on its knowledge base to provide factual information, but it did not account for health-related changes over time (Ouyang et al., 2025). Some studies combined an LLM with online web searches to identify evidence-based misinformation (Tian et al., 2024; Li et al., 2024a; Dey et al., 2024), which mitigated the limitations of static knowledge base. These systems enhanced evidence retrieval by incorporating information from diverse online sources (Shi et al., 2025). However, web searches introduced new challenges like retrieving low-quality or misleading sources. To address these issues, our study integrated both static and dynamic retrieved evidence to ensure the availability of relevant, real-time information for countering misinformation.
@@ 2.2 LLM Agents for Social Response Generation
Integrating LLMs across multiple agents has dramatically improved response generation (Zhang et al., 2024; Zhao et al., 2024). Multi-agent LLM systems featured several specialized agents working together to tackle complex tasks, utilizing LLMs to boost the quality and efficiency of interactions (Yang et al., 2025; Guo et al., 2024; Cruz, 2024). Many studies utilized multiple LLM-based multi-agent architectures to improve information retrieval and response generation (Salve et al., 2024), tackling tasks that required long-context understanding (Zhang et al., 2024), and enhancing recommendation quality and user engagement (Fang et al., 2024). For example, Clarke et al. (2022) implemented a framework where agents specialized in different tasks but coordinated to deliver cohesive responses, utilizing a central management system. These multi-agent approaches demonstrated the effectiveness of breaking down complex response generation into modular, specialized components. Inspired by this, our Multi-agent Retrieval-Augmented Framework employed sequentially organized agents, each with distinct capabilities, to retrieve, filter, summarize, generate, and refine evidence-based counterspeech.
Figure 1: Overview of our proposed Multi-Agent Retrieval-Augmented Framework for addressing health misinformation. First, the Static Retrieval Agent collects evidence from a local database, while the Dynamic Retrieval Agent gathers real-time online information. Next, the Summarization Agent filters and condenses the retrieved evidence for clarity. The Counterspeech Agent generates a response, and the Refinement Agent refines it to ensure clarity, respectfulness, and strong evidential support. This process ensures effective, evidence-based counterspeech.
@@ 3.1 Generation
The counterspeech generation process involves leveraging LLMs to create effective, informative, and polite responses that counter health misinformation. The primary objective is to generate responses that not only provide factual corrections but also address misinformation in a respectful and engaging manner. Our proposed multi-agent system uses specialized agents that work in sequence to retrieve, filter, and process evidence, thereby optimizing the RAG pipeline (Figure 1).
Information Retrieval To incorporate static and dynamic evidence, two separate agents are employed to retrieve evidence: one is a Static Retrieval Agent and another one is a Dynamic Retrieval Agent. These Agents collect information based on the health misinformation posts. The Static Retrieval Agent conducts an inquiry within a structured local knowledge base by employing a dual approach that integrates vector-based semantic search with keyword-based exact matching (Yuan et al., 2024; Sawarkar et al., 2024), ensuring that only valid medical knowledge is included in this process. In parallel, the Dynamic Retrieval Agent performs real-time web searches using the DuckDuckGo API, introducing the latest health misinformation corrections from authoritative sources. The two sources of knowledge are combined to produce the final evidence.
@@ 3.2 Evaluation
We employ a comprehensive evaluation framework based on four key criteria: politeness, relevance, informativeness, and factual accuracy. These criteria ensure that the responses are engaging, respectful, factually sound, and effective in addressing health misinformation.
@@
(b) Relevance: Relevant responses ensure the response directly addresses the specific health misinformation claim (Yue et al., 2024; He et al., 2023). We assess relevance using the BM25 relevance scoring algorithm (Robertson et al., 2009), which measures how well the generated response aligns with the original health misinformation content.
(c) Informativeness: Informativeness reflects the amount of useful, relevant, and evidence-based content shared in the response. A highly informative counterspeech response not only corrects health misinformation but also provides actionable insights and explanations that improve user understanding and support informed decision-making (Chung et al., 2024). To measure informativeness, we employ LLM as an evaluator (Li et al., 2024b) with detailed instructions, assessing if the response provides valuable information (Hong et al., 2025).
(d) Factual Accuracy: Factual accuracy demonstrates the reliability and trustworthiness of the generated response by ensuring that the information is correct and supported by evidence (Zhou et al., 2024). In counterspeech generation, presenting scientifically accurate facts can effectively correct misinformation and maintain user trust (Yue et al., 2024). We employ LLM evaluation (Li et al., 2024b) to check how accurate the generated answers are by comparing them to the gathered evidence. The model rates how accurate a response is by checking if it correctly presents scientific facts and does not include any false information.
@@ 4.1 Dataset
We utilized PRW API  [^6] https://praw.readthedocs.io/ to collect health misinformation from Reddit, focusing on topics related to the coronavirus disease (COVID-19), influenza (Flu), and human immunodeficiency virus (HIV), as these topics have been the center of conspiracy theories and misinformation in recent times. We collected Reddit posts and comments containing health-related keywords (e.g., ”vaccines,” ”COVID-19,” ”alternative medicine”, etc. See details in Appendix A.1.2). Finally, we obtained about 4,968 posts and 17,223 comments from high-engagement subreddits (e.g., r/health or r/science, etc. See details in Appendix A.1.1).
After collecting the Reddit dataset, we employed human annotators to identify and filter health misinformation (Song et al., 2025b). Considering the annotation cost, we only extracted 1000 samples for annotation. Five annotators with backgrounds in information science were recruited for the annotation. To ensure consistency and reliability in annotations, all annotators were provided with comprehensive and detailed annotation guidelines Appendix A.2.1. We finally obtained 330 posts containing health misinformation.
To further enrich our dataset, we developed an automatic classifier trained on the annotated data, enabling the efficient collection of additional health misinformation posts from unannotated data. We employed the RoBERTa-large model (Liu et al., 2019), trained on our manually annotated dataset, which achieves a moderate accuracy with an F1 score of 0.76 (Implementation details in Appendix A.2.2). We applied it to the remaining dataset, identifying 831 additional posts containing health misinformation. To validate the reliability of the classifier, we extracted 100 samples to conduct human validation (See details in Appendix A.2.3). The final dataset contained 1,161 posts labeled as health misinformation.